### v1

# The Waterbed Effect on Quasiperiodic Disturbance Observer: Avoidance of Sensitivity Tradeoff with Time Delays

Abstract
In linear time-invariant systems, the sensitivity function to disturbances is designed under a sensitivity tradeoff known as the waterbed effect. To compensate for a quasiperiodic disturbance, a quasiperiodic disturbance observer using time delays was proposed. Its sensitivity function avoids the sensitivity tradeoff, achieving wideband harmonic suppression without amplifying aperiodic disturbances or shifting harmonic suppression frequencies. However, its open-loop transfer function is not rational and does not satisfy the assumptions of existing Bode sensitivity integrals due to its time delays. This paper provides Bode-like sensitivity integrals for the quasiperiodic disturbance observer in both continuous-time and discrete-time representations and clarifies the avoided sensitivity tradeoff with time delays.
keywords:
Disturbance observer, time delays, harmonics, repetitive control, iterative learning control, sensitivity function, waterbed effect, Bode sensitivity integral
thanks: This work was supported by JSPS KAKENHI Grant Number JP25K17560.
1 Introduction
One of the fundamental issues in automatic control is disturbance compensation. In practical systems, disturbances always exist and impair the precision of automatic control. For linear time-invariant systems, we design a sensitivity function $S(s)\coloneqq 1/(1+\Gamma(s))$ based on an open-loop transfer function $\Gamma(s)$, while tradeoffs restrict the design. One tradeoff lies between the sensitivity function $S(s)\coloneqq 1/(1+\Gamma(s))$ and complementary sensitivity function $T(s)\coloneqq\Gamma(s)/(1+\Gamma(s))$, which are restricted by the unit sum $S(s)+T(s)=1$. The sensitivity and complementary sensitivity functions represent disturbance-suppression performance and robust stability against modeling errors, respectively. It is necessary to reduce their gains appropriately to improve performance under the unit-sum tradeoff. For the sensitivity function, there exists another tradeoff of the sensitivity integral given by Bode (1945). This Bode sensitivity integral (Freudenberg and Looze (1985); Ding et al. (2022); Nagarsheth and Sharma (2021)) states that a stable sensitivity function satisfies
$\int_{0}^{\infty}\ln|S(j\omega)|d\omega=\pi\sum_{i}\mathrm{Re}[q_{i}]-\frac{\pi}{2}\lim_{s\to\infty}s\Gamma(s),$ (1)
where $q_{i}$ are the unstable right-half-plane poles of a proper rational function $\Gamma(s)$. Also, Wu and Jonckheere (1992); SUNG and HARA (1988); Mohtadi (1990); Emami-Naeini and de Roover (2019) provide the Bode sensitivity integral for discrete-time control. Based on a rational open-loop transfer function $\tilde{\Gamma}(z)$, the stable sensitivity function $\tilde{S}(z)\coloneqq 1/(1+\tilde{\Gamma}(z))$ satisfies
$\int_{0}^{2\pi}\ln|\tilde{S}(e^{j\omega})|d\omega$ $=2\pi\sum_{i}\ln|\tilde{q}_{i}|$
$\hskip 10.00002pt-2\pi\ln|\lim_{z\to\infty}\tilde{\Gamma}(z)+1|,$ (2)
where $\tilde{q}_{i}$ are unstable poles outside the closed-unit disk of a proper rational function $\tilde{\Gamma}(z)$. The Bode sensitivity integral is called the waterbed effect (Skogestad and Postlethwaite (2005); Costa-Castelló and Dormido (2015)), as the gain increases at some frequencies while decreasing at other frequencies. However, these Bode sensitivity integrals are applicable only to proper rational open-loop transfer functions.
Periodic disturbances often occur in automatic mechanical and electrical systems. To compensate for harmonics of a periodic disturbance, time delays are essential. Repetitive control (Hara et al. (1988)), iterative learning control (Wang et al. (2009); Bristow et al. (2006)), and periodic disturbance observers (Muramatsu and Katsura (2018)) have been proposed using time delays for an internal model of a periodic disturbance. However, in designing the sensitivity function with time delays, a waterbed-effect-like tradeoff exists between the wideband harmonic suppression, amplification of aperiodic disturbances, and deviation of harmonic suppression frequencies (Pipeleers et al. (2008); Chen and Tomizuka (2014); Nie et al. (2021); Tanaka and Muramatsu (2023); Yang et al. (2023); Lai et al. (2021); Li et al. (2023)). To overcome the tradeoff, Muramatsu (2025) proposed a quasiperiodic disturbance observer (QDOB) using time delays, which achieved wideband harmonic suppression, non-amplification of aperiodic disturbances, and non-deviation of harmonic suppression frequencies, simultaneously. From the perspective of the Bode sensitivity integral, this paper examines the waterbed effect on the QDOB in both continuous-time and discrete-time representations.
The Bode sensitivity integral for disturbance observers based on rational transfer functions has been studied by Sariyildiz and Ohnishi (2013a, b). However, the results cannot be applied to the QDOB by Muramatsu (2025), whose open-loop transfer function is
$\Gamma(s)$ $=\frac{\omega_{\mathrm{c}}L}{2}\frac{1+\Phi(s)}{1-\Phi(s)}\frac{\omega_{\mathrm{b}}}{s+\omega_{\mathrm{b}}},$ (3)
where $\omega_{\mathrm{b}},\omega_{\mathrm{c}},L\in\mathbb{R}$. The function $\Phi(s)$ is a linear-phase low-pass filter that can be expressed as
$\Phi(s)=e^{-\kappa Ts}\prod_{i=1}^{l}\sum_{n=-N}^{N}\frac{\alpha_{i}(n)}{\gamma_{i}}e^{(n-N)T\bar{U}_{i}s},$ (4)
where $\kappa\coloneqq\bar{L}-N\sum_{i=1}^{l}\bar{U}_{i}\geq 1$, $\bar{U}_{i}\coloneqq\mathrm{round}(U_{i}/T)$, and $\bar{L}\coloneqq\mathrm{round}(L/T)$. The open-loop transfer function has an infinite number of zeros by $1+\Phi(s)$ and an infinite number of poles by $1-\Phi(s)$. Consequently, the open-loop transfer function $\Gamma(s)$ is not rational and is a meromorphic function with time delays.
Several studies have extended the class of open-loop transfer functions applicable to the Bode sensitivity integral. Freudenberg and Looze (1987) investigated the Bode sensitivity integral with an open-loop transfer function $\Gamma_{0}(s)e^{-\tau s}$ combining a rational function $\Gamma_{0}(s)$ and time delay $e^{-\tau s}$. Chang et al. (2023) and Songchon and Longman (2000) investigated the integrals with fractional-order proportional-integral-derivative control and discrete-time iteration learning control, respectively. Furthermore, Gomez and Goodwin (1998) obtained the Bode sensitivity integral for meromorphic functions with time delays; however, it cannot address an infinite number of zeros of the sensitivity function on the imaginary axis, such as the QDOB.
The main results of this paper are theorems of Bode-like sensitivity integrals for the QDOB in both continuous-time and discrete-time representations. Through the theorems and numerical integration results, this paper demonstrates how the QDOB avoids the waterbed effect.
2 Preliminaries
This section presents the QDOB proposed by Muramatsu (2025) as preliminaries. Suppose the single-input-single-output plant
$\mathcal{L}[y(t)]=P(s)\mathcal{L}[u(t)+v(t)],$ (5)
where
$P(s)$ $\coloneqq(1+\Delta(s))P_{\mathrm{n}}(s)$ (6a)
$P_{\mathrm{n}}(s)$ $\coloneqq\frac{b_{n}s^{n}+b_{n-1}s^{n-1}+\cdots+b_{1}s+b_{0}}{s^{m}+a_{m-1}s^{m-1}+\cdots+a_{1}s+a_{0}}$ (6b)
$\Delta(s)$ $\coloneqq\frac{\beta_{g}s^{g}+\beta_{g-1}s^{g-1}+\cdots+\beta_{1}s+\beta_{0}}{s^{h}+\alpha_{h-1}s^{h-1}+\cdots+\alpha_{1}s+\alpha_{0}}e^{-\gamma s}$ (6c)
with $g,h\in\mathbb{Z}_{\geq 0}, n<mn<m,$ and $\gamma\in\mathbb{R}_{\geq 0}$. The symbols $t\in\mathbb{R}$, $u(t)\in\mathbb{R}$, $v(t)\in\mathbb{R}$, $y(t)\in\mathbb{R}$, $P(s)\in\mathbb{C}$, $P_{\mathrm{n}}(s)\in\mathbb{C}$, $\Delta(s)\in\mathbb{C}$, and $\mathcal{L}$ denote the time, input, exogenous signal, output, plant, nominal plant, stable modeling error, and Laplace transform operator, respectively. Assume that the strictly proper transfer function $P_{\mathrm{n}}(s)$ has no zeros or poles in the closed right-half plane. For the plant, the disturbance $d(t)$ including the exogenous signal and effects of the modeling error is defined as
$\mathcal{L}[d(t)]\coloneqq\mathcal{L}[v(t)]+\Delta(s)\mathcal{L}[u(t)+v(t)],$ (7)
and the plant is rewritten as
$\mathcal{L}[y(t)]=P_{\mathrm{n}}(s)\mathcal{L}[u(t)+d(t)].$ (8)
Using the lifting transform $\mathcal{G}$ and the discrete-time Fourier transform $\mathcal{F}$:
$\mathcal{G}[d]_{L,\tau}(c)\coloneqq d(cL+\tau),\ L\in\mathbb{R}_{>0},\ \tau\in[0,L),\ c\in\mathbb{Z}$ (9a)
$\mathcal{F}[D](\Omega)\coloneqq\sum_{c=-\infty}^{\infty}D(c)e^{-j\Omega c},$ (9b)
where $\Omega\in[0,2\pi)$ [rad/sample] denotes the normalized angular frequency. Then, Muramatsu (2025)’s definition of the quasiperiodic disturbance can be stated as Definition 1.
Definition 1
A disturbance $d\in\mathcal{D}_{L}$ is called quasiperiodic with respect to a period $L\in\mathbb{R}_{>0}$ and separation frequency $\rho\in[0,\pi/L)$ if $d\in\mathcal{P}_{L,\rho}\subset\mathcal{D}_{L}$.
$\mathcal{D}_{L}$ $\coloneqq\{d:\mathbb{R}\to\mathbb{R}|\mathcal{G}[d]_{L,\tau}\in\ell^{1}(\mathbb{Z}),\ \forall\tau\in[0,L)\}$ (10a)
$\mathcal{P}_{L,\rho}$ $\coloneqq\{d\in\mathcal{D}_{L}|\mathcal{F}[\mathcal{G}[d]_{L,\tau}](\Omega)=0,$
$\hskip 60.00009pt\forall\tau\in[0,L),\ \forall\Omega\in(\rho L,\pi]\}.$ (10b)
This definition is based on Muramatsu (2022)’s quasiperiodicity. Let $\omega_{0}\coloneqq 2\pi/L$ and $n\omega_{0}$ be the fundamental and harmonic frequencies of the quasiperiodic disturbance, respectively.
The QDOB, which estimates and compensates for the quasiperiodic disturbance, is
$\mathcal{L}[\hat{d}(t)]=Q(s)\mathcal{L}[\xi(t)-u(t)]$ (11a)
$\mathcal{L}[\xi(t)]=B(s)P_{\mathrm{n}}^{-1}(s)\mathcal{L}[y(t)]$ (11b)
$u(t)=r(t)-\hat{d}(t),$ (11c)
where $\hat{d}(t)\in\mathbb{R}$ is the estimated quasiperiodic disturbance. The QDOB is composed of a first-order low-pass filter
$B(s)\coloneqq\frac{\omega_{\mathrm{b}}}{s+\omega_{\mathrm{b}}}$ (12)
and the Q-filter
$Q(s)$ $\coloneqq\frac{\omega_{\mathrm{c}}L(1+\Phi(s))}{(\omega_{\mathrm{c}}L+2)+(\omega_{\mathrm{c}}L-2)\Phi(s)}$ (13a)
$\omega_{\mathrm{c}}$ $\coloneqq\frac{2}{L}\tan{\left(\frac{L}{2}\rho\right)},$ (13b)
using the linear-phase low-pass filter $\Phi(s)$:
$\Phi(s)\coloneqq e^{-\kappa Ts}\prod_{i=1}^{l}\varphi_{i}(s)$ (14a)
$\varphi_{i}(s)\coloneqq\frac{1}{\gamma_{i}}\sum_{n=-N}^{N}\alpha_{i}(n)e^{(n-N)T\bar{U}_{i}s}$ (14b)
$\alpha_{i}(n)\coloneqq w(n,N)h(n,\omega_{i},U_{i})$ (14c)
[…]
$\gamma_{i}\coloneqq\max_{\Omega\in[0,\pi]}\bigg\{\alpha_{i}(0)+2\sum_{n=1}^{N}\alpha_{i}(n)\cos(n\Omega)\bigg\}$ (14k)
$U_{i}\coloneqq\left\{\begin{array}[]{cl}T&\mathrm{if}\ i=1\\ {\pi}/{\omega_{i-1}}&\mathrm{otherwise}\end{array}\right$. (14n)
$\omega_{i}\coloneqq\left\{\begin{array}[]{cl}\omega_{\mathrm{a}}&\mathrm{if}\ i=l\\ {2c\pi}/{U_{i}}&\mathrm{otherwise},\end{array}\right.\ c=\frac{1}{2}\left(\frac{T\omega_{\mathrm{a}}}{\pi}\right)^{1/l}$ (14q)
$N\coloneqq\min\{\max(\mathcal{N}),\ N_{\mathrm{max}}\}$ (14r)
$\mathcal{N}\coloneqq\{n\in\mathbb{Z}_{>0}|n\leq(L-T)/\textstyle\sum_{i=1}^{l}U_{i}\}.$ (14s)
The parameters $T\in\mathbb{R}_{>0}$, $\omega_{\mathrm{a}}\in\mathbb{R}_{>0}$, $l\in\mathbb{Z}_{>0}$, and $N_{\mathrm{max}}\in\mathbb{Z}_{>0}$ denote the sampling time, cutoff frequency, number of stages, and maximum order, respectively. This linear-phase low-pass filter can be equivalently expressed as (4). The Q-filter is based on the periodic/aperiodic separation filter (Muramatsu (2022); Muramatsu and Katsura (2019)) whose time delays are multiplied by zero-phase low-pass filters. Consequently, the open-loop transfer function $\Gamma(s)$ of the QDOB is
$\Gamma(s)$ $=\frac{\omega_{\mathrm{c}}L}{2}\frac{1+\Phi(s)}{1-\Phi(s)}\frac{\omega_{\mathrm{b}}}{s+\omega_{\mathrm{b}}},$ (0)
and its sensitivity and complementary sensitivity functions are
$S(s)$ $\coloneqq\frac{1}{1+\Gamma(s)},\ T(s)\coloneqq\frac{\Gamma(s)}{1+\Gamma(s)},$ (15)
respectively.
The QDOB is discretized by two methods. The Q-filter $Q(s)$ is discretized by $e^{-Ts}\to z^{-1}$, and the transfer function $B(s)P_{\mathrm{n}}^{-1}(s)$ is discretized by the backward Euler method $s\to(1-z^{-1})/T$. Then, the open-loop transfer function in the discrete-time representation is
$\tilde{\Gamma}(z)$ $=\frac{\omega_{\mathrm{c}}L}{2}\frac{1+\tilde{\Phi}(z)}{1-\tilde{\Phi}(z)}\frac{\omega_{\mathrm{b}}Tz}{(1+\omega_{\mathrm{b}}T)z-1},$ (16)
where
$\tilde{\Phi}(z)\coloneqq z^{-\kappa}\prod_{i=1}^{l}\tilde{\varphi}_{i}(z)$ (17a)
$\tilde{\varphi}_{i}(z)\coloneqq\frac{\sum_{n=-N}^{N}w(n,N)h(n,\omega_{i},U_{i})z^{(n-N)\bar{U}_{i}}}{\sum_{n=-N}^{N}w(n,N)h(n,\omega_{i},U_{i})}.$ (17b)
Accordingly, the sensitivity and complementary sensitivity functions in the discrete-time representation are
$\tilde{S}(z)$ $\coloneqq\frac{1}{1+\tilde{\Gamma}(z)},\ \tilde{T}(z)\coloneqq\frac{\tilde{\Gamma}(z)}{1+\tilde{\Gamma}(z)},$ (18)
respectively.
(a) Sensitivity function $\tilde{S}(e^{j\Omega})$.
(b) Complementary sensitivity function $\tilde{T}(e^{j\Omega})$.
(c) Open-loop transfer function $\tilde{\Gamma}(e^{j\Omega})$.
Figure 1: Bode plots of the QDOB in the discrete-time representation with $l=3$, $N_{\mathrm{max}}=256$, $\omega_{0}=1$ rad/s, $\omega_{\mathrm{a}}=10$ rad/s, $\omega_{\mathrm{b}}=100$ rad/s, and $T=1$ ms.
Fig. 1 shows the Bode plots of the sensitivity, complementary sensitivity, and open-loop transfer functions in the discrete-time representation. As shown in Fig. 1(a), the QDOB can expand the harmonic suppression bandwidth by increasing the separation frequency $\rho$ without the waterbed effect. See Muramatsu (2025) for the detailed derivation, design, analyses, and discretization of the QDOB, including the Bode plots of the functions in the continuous-time representation.
Remark 2 describes the modifications to Muramatsu (2025), and Remark 3 introduces how to design the sensitivity function for the QDOB on the basis of Muramatsu (2025).
Remark 2
The normalization $1/\gamma_{i}$ of (14b) for the linear-phase low-pass filter $\Phi(s)$ has been modified so that the gain of the filter satisfies $|\Phi(j\omega)|\leq 1$. In Muramatsu (2025), it was normalized by the DC gain $\gamma_{i}=\alpha_{i}(0)+2\sum_{n=1}^{N}\alpha_{i}(n)$ by neglecting passband ripple of the filter. Additionally, rounded $\bar{U}_{i}T$ and $\bar{L}T$ are used for the continuous-time representation instead of $U_{i}$ and $L$ in the time delays. Lastly, minor errors in the set $(\rho L,\pi]$ for $\Omega$ of (10) in Definition 1 and of the upper limit of the summation of (14s) have been corrected.
Remark 3
The crucial hyperparameters of the QDOB for the sensitivity design are $\omega_{\mathrm{a}}$, $\omega_{\mathrm{b}}$, and $\rho$. The sensitivity function suppresses harmonics at frequencies below $\omega_{\mathrm{a}}$, and its gain becomes less than 1 at frequencies below $\omega_{\mathrm{b}}$. Additionally, the complementary sensitivity function decreases the gain for robust stability at frequencies above $\omega_{\mathrm{b}}$. The angular frequencies $\omega_{\mathrm{a}}$ and $\omega_{\mathrm{b}}$ are adjusted to design the sensitivity and complementary sensitivity functions within the condition $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$, which maintains the phase of the open-loop transfer function within 90-90 deg. to $90$ deg. at frequencies below the Nyquist frequency. The separation frequency $\rho\in[0,\pi/L)$ changes the harmonic suppression bandwidth such that $20\log|S(j\omega)|<-3$ dB $\forall\omega\in[n\omega_{0}-\rho,n\omega_{0}+\rho]\cap[0,\omega_{\mathrm{a}}]$ around the harmonic frequencies $n\omega_{0}$.
Based on the preliminaries, this paper discusses the waterbed effect on the sensitivity functions $S(s)$ and $\tilde{S}(z)$ in the continuous-time and discrete-time representations.
3 Waterbed Effect on QDOB
3.1 Main Results
The main results of this paper are two theorems that apply a Bode-like sensitivity integral to the QDOB in the continuous-time or discrete-time representation. Two lemmas are provided for the proofs of the theorems.
Lemma 4
Suppose (14) and $g\in\mathbb{R}_{\geq 0}$; then,
$\mathcal{S}_{g}\coloneqq\{s\in\mathbb{C}_{+}\mid\mathrm{Re}[s]>g\}$ (19a)
$|\Phi(s)|<\max_{\mathrm{Re}[s]=g}|\Phi(s)|,\ \forall s\in\mathcal{S}_{g}.$ (19b)
{pf}
The proof is in Appendix A.
Lemma 5
The open-loop transfer function (3) has no poles in the open right-half plane.
{pf}
The proof is in Appendix B.
Theorem 6
Suppose the open-loop transfer function (3). Then, the sensitivity function (15) in the continuous-time representation satisfies
$\lim_{\varepsilon\to 0^{+}}\int_{0}^{\infty}\ln|S(\varepsilon+j\omega)|d\omega=-\frac{\pi\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}{4}.$ (20)
{pf}
The proof is in Appendix C.
Theorem 7
Suppose the open-loop transfer function (16). Then, the sensitivity function (18) in the discrete-time representation satisfies
$\lim_{\varepsilon\to 1^{+}}\int_{0}^{2\pi}\ln|\tilde{S}(\varepsilon e^{j\Omega})|d\Omega$
$=2\pi\ln(2+2\omega_{\mathrm{b}}T)-2\pi\ln(2+2\omega_{\mathrm{b}}T+\omega_{\mathrm{b}}\omega_{\mathrm{c}}LT).$ (21)
{pf}
The proof is in Appendix D.
From the theorems, the increase in $\omega_{\mathrm{b}}$ and/or $\rho$ reduces the gain of the sensitivity function as described in Remark 3, while the Bode-like sensitivity integrals (20) and (7) decrease in both continuous-time and discrete-time representations. Note that $\omega_{\mathrm{c}}$ is determined by $\rho$ according to (13b), and they are in a monotonic relationship. This suggests that the QDOB can extend the harmonic suppression bandwidth without a sensitivity tradeoff. In contrast, $\omega_{\mathrm{a}}$ does not affect the integrals. This implies that the band stops around harmonic frequencies below $\omega_{\mathrm{a}}$ are restricted by the sensitivity tradeoff. Fig. 2 illustrates the relationship among the gain of the sensitivity function, $\omega_{\mathrm{a}}$, $\omega_{\mathrm{b}}$, $\omega_{\mathrm{c}}$, and $\rho$. It shows that $\omega_{\mathrm{b}}$ extends the suppression bandwidth horizontally, and $\omega_{\mathrm{c}}L$ deepens the attenuation depth vertically. Furthermore, the waterbed effect on $\omega_{\mathrm{a}}$ is represented as the red region of the reduced gain and the blue region of the increased gain, while the gain does not exceed 0 dB.
Figure 2: Illustrative diagram of the gain of the sensitivity function for the QDOB.
Theorems 6 and 7 employ the Bode-like sensitivity integrals based on the limits avoiding the singularities on the imaginary axis and unit disk because the sensitivity function has an infinite and finite number of roots, respectively. Although the QDOB does not satisfy the assumptions of the existing Bode sensitivity integrals, the Bode-like sensitivity integrals (20) and (7) turn out to be consistent with the Bode sensitivity integrals (1) and (1), respectively.
Table 1: Parameter settings and theoretical values of the Bode-like sensitivity integrals with $T=1$ ms.
$\delta_{\mathrm{ct}}$ $\delta_{\mathrm{dt}}$ $\omega_{0}$ $\omega_{\mathrm{b}}$ $\rho$
$\mathrm{P}_{1}$ -51.0 -0.18 1 rad/s 100 rad/s 0.1 rad/s
$\mathrm{P}_{2}$ -157.1 -0.55 1 rad/s 100 rad/s 0.25 rad/s
$\mathrm{P}_{3}$ -314.2 -0.84 10 rad/s 400 rad/s 1.476 rad/s
$\mathrm{P}_{4}$ -314.2 -0.97 1 rad/s 200 rad/s 0.25 rad/s
[…]
3.2 Numerical Results
Numerical examples are provided to verify the convergence of the integrals: $\delta_{\mathrm{ct}}\coloneqq\int_{0}^{w}\ln|S(j\omega)|d\omega$ and $\delta_{\mathrm{dt}}\coloneqq\int_{0}^{w}\ln|\tilde{S}(e^{j\Omega})|d\Omega$ with respect to the upper limit of integration $w$ toward (20) and (7). For validation, parameter settings of Table 1 for the hyperparameters: $\omega_{0}$, $\omega_{\mathrm{b}}$, and $\rho$ are considered. Accordingly, (20) and (7) are expressed by them as
$\delta_{\mathrm{ct}}$ $=-\frac{\pi\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}{4}=-\frac{\pi\omega_{\mathrm{b}}}{2}\tan{\left(\frac{\pi\rho}{\omega_{0}}\right)}$ (22a)
$\delta_{\mathrm{dt}}$ $=2\pi\ln(2+2\omega_{\mathrm{b}}T)-2\pi\ln(2+2\omega_{\mathrm{b}}T+\omega_{\mathrm{b}}\omega_{\mathrm{c}}LT)$
$=2\pi\ln(1+\omega_{\mathrm{b}}T)$
$\hskip 20.00003pt-2\pi\ln\left(1+\omega_{\mathrm{b}}T+\omega_{\mathrm{b}}T\tan{(\pi\rho/\omega_{0})}\right).$ (22b)
(a) $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$
(b) $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$
(c) $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$
(d) $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$
Figure 3: Convergence of the Bode-like sensitivity integrals toward the theoretical values with $l=3$, $N_{\mathrm{max}}=256$, $\omega_{\mathrm{a}}=10$ rad/s, $T=1$ ms, and Table 1.
The numerical results are shown in Fig. 3. In the continuous-time cases of Fig. 3(a) and (c), the integral values asymptotically converge toward $\delta_{\mathrm{ct}}$ as $w$ increases. In the discrete-time cases of Fig. 3(b) and (d), the integral values converge to $\delta_{\mathrm{dt}}$ at $w=\pi$. Note that
$\int_{0}^{2\pi}\ln|\tilde{S}(e^{j\Omega})|d\Omega=2\int_{0}^{\pi}\ln|\tilde{S}(e^{j\Omega})|d\Omega$ (23)
based on $|\tilde{S}(e^{j\Omega})|=|\tilde{S}(e^{-j\Omega})|$ and $|\tilde{S}(e^{j\Omega})|=|\tilde{S}(e^{j(\Omega+2\pi)})|$.
Although Theorems 6 and 7 hold regardless of the relationship between $\omega_{\mathrm{a}}$ and $\omega_{\mathrm{b}}$, the convergences of the integral values are different between $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$ [Fig. 3(a) and (b)] and $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$ [Fig. 3(c) and (d)]. If $\omega_{\mathrm{a}}>\omega_{\mathrm{b}}$, the gains $|S(j\omega)|$ and $|\tilde{S}(e^{j\Omega})|$ repeatedly increase and decrease while converging. This increase corresponds to the amplification of aperiodic disturbances and can cause a shift in harmonic suppression frequencies, which usually happen with conventional repetitive control. Thus, the avoidance of the sensitivity tradeoff in Theorems 6 and 7 does not prevent local amplification of the gain.
To avoid the local amplification, the phase of the open-loop transfer function within $\pm 90$ deg. is essential. To discuss this, this paper considers the following proposition.
Proposition 8
Suppose an open-loop transfer function $\Gamma(j\omega)$ and sensitivity function $S(j\omega)=(1+\Gamma(j\omega))^{-1}$. Then, $|\angle\Gamma(j\omega)|\leq\pi/2\Rightarrow|S(j\omega)|\leq 1$.
{pf}
$|\angle\Gamma(j\omega)|\leq\pi/2\Rightarrow\mathrm{Re}[\Gamma(j\omega)]\geq 0\Rightarrow(1+\mathrm{Re}[\Gamma(j\omega)])^{2}+\mathrm{Im}[\Gamma(j\omega)]^{2}\geq 1\Rightarrow|S(j\omega)|=((1+\mathrm{Re}[\Gamma(j\omega)])^{2}+\mathrm{Im}[\Gamma(j\omega)]^{2})^{-1/2}\leq 1$. On the basis of Remark 3 and Proposition 8, the parameter design $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$, which makes the phase of the open-loop transfer function within $\pm 90$ deg. at frequencies below the Nyquist frequency, prevents local amplification of the gain at the frequencies. Furthermore, the separation frequency $\rho$ does not affect the phase of the open-loop transfer function $\Gamma(s)$, and the phase exists within $\pm 90$ deg. regardless of $\rho$, as shown in Fig. 1(c). This indicates that the gain of the sensitivity function always satisfies $|S(j\omega)|\leq 1$ regardless of $\rho$ if $\omega_{\mathrm{a}}\ll\omega_{\mathrm{b}}$. These results explain how the increase in the separation frequency $\rho$ extends the harmonic suppression bandwidth while decreasing the Bode-like sensitivity integral, without causing amplification at other frequencies below the Nyquist frequency. For the continuous-time representation, at frequencies beyond the Nyquist frequency, the low-pass filter $B(j\omega)$ and sensitivity function $S(j\omega)$ asymptotically approach zero and one, respectively, as the frequency increases.
4 Conclusion
This paper examined the waterbed effect of the QDOB proposed by Muramatsu (2025), which estimates and compensates for the quasiperiodic disturbance of Definition 1. Compared to conventional periodic-disturbance-compensation control, the advantage of the QDOB is the simultaneous realization of the wideband harmonic suppression, non-amplification of aperiodic disturbances, and proper harmonic suppression frequencies.
In this paper, Theorems 6 and 7 indicated the avoidance of the sensitivity tradeoff, where the extension of harmonic suppression bandwidth for the sensitivity function also reduces the Bode-like sensitivity integral. Moreover, the numerical results and Proposition 8 showed the waterbed effect on the QDOB, where $|S(j\omega)|\leq 1$ achieved by the phase of the open-loop transfer function within $\pm 90$ deg. attains the non-amplification of aperiodic disturbances and proper harmonic suppression frequencies.
Appendix A Proof of Lemma 0
The linear-phase low-pass filter $\Phi(s)$ in (14) is analytic and can be expressed as (4). Let
$\mathcal{E}_{g}\coloneqq\{\xi\in\mathbb{C}\mid|\xi|<e^{-Tg}\},\ \xi\coloneqq e^{-Ts}$ (24a)
$H(\xi)\coloneqq\prod_{i=1}^{l}\sum_{n=-N}^{N}\frac{\alpha_{i}(n)}{\gamma_{i}}\xi^{(N-n)\bar{U}_{i}}.$ (24b)
Based on (4),
$\xi\in\mathcal{E}_{g}\Leftrightarrow s\in\mathcal{S}_{g},\ |\xi|=e^{-Tg}\Leftrightarrow\mathrm{Re}[s]=g$ (25a)
$\Phi(s)=e^{-\kappa Ts}H(e^{-Ts}).$ (25b)
Since $H(\xi)$ is nonconstant and analytic in the open disk $\mathcal{E}_{g}$ and continuous on the closure of $\mathcal{E}_{g}$, the maximum modulus principle gives
$|H(\xi)|<\max_{|\xi|=e^{-Tg}}|H(\xi)|,\ \forall\xi\in\mathcal{E}_{g}.$ (26)
By the change of variable $\xi=e^{-Ts}$,
$|H(e^{-Ts})|=|e^{\kappa Ts}||\Phi(s)|=e^{\kappa T\mathrm{Re}[s]}|\Phi(s)|$ (27a)
$\max_{|\xi|=e^{-Tg}}|H(\xi)|=\max_{\mathrm{Re}[s]=g}|H(e^{-Ts})|$
$\hskip 20.00003pt=\max_{\mathrm{Re}[s]=g}|e^{\kappa Ts}\Phi(s)|=e^{\kappa Tg}\max_{\mathrm{Re}[s]=g}|\Phi(s)|.$ (27b)
Hence,
$e^{\kappa T\mathrm{Re}[s]}|\Phi(s)|<e^{\kappa Tg}\max_{\mathrm{Re}[s]=g}|\Phi(s)|,\ \forall s\in\mathcal{S}_{g}.$ (28)
Because $e^{\kappa T\mathrm{Re}[s]}>e^{\kappa Tg}\geq 1$, $\forall s\in\mathcal{S}_{g}$, we have
$|\Phi(s)|<\max_{\mathrm{Re}[s]=g}|\Phi(s)|,\ \forall s\in\mathcal{S}_{g}.$ (29)
Appendix B Proof of Lemma 0
The poles of (3) are roots of $s+\omega_{\mathrm{b}}$ and $1-\Phi(s)$. The root of $s+\omega_{\mathrm{b}}$ is $s=-\omega_{\mathrm{b}}$ in the open left-half plane. The linear-phase low-pass filter $\Phi(s)$ in (14) consists of multiple linear-phase low-pass filters $\varphi_{i}(s)$ in (14b). Its gain is
$|\varphi_{i}(j\omega)|$ $=\frac{1}{\gamma_{i}}\left(\alpha_{i}(0)+\sum_{n=1}^{N}\alpha_{i}(n)(e^{jnT\bar{U}_{i}\omega}+e^{-jnT\bar{U}_{i}\omega})\right)$
$=\frac{1}{\gamma_{i}}\left(\alpha_{i}(0)+2\sum_{n=1}^{N}\alpha_{i}(n)\cos(nT\bar{U}_{i}\omega)\right)$ (30)
using the property $\alpha_{i}(n)=\alpha_{i}(-n)$ of even functions. Eqs. (14b) and (14k) yield $|\varphi_{i}(j\omega)|\leq 1$, $\forall\omega\in\mathbb{R}$, and
$\left|\Phi(j\omega)\right|=\prod_{i=1}^{l}\left|\varphi_{i}(j\omega)\right|\leq 1,\ \forall\omega\in\mathbb{R}.$ (31)
Lemma 4 gives
$|\Phi(s)|<\max_{\mathrm{Re}[s]=0}|\Phi(s)|,\ \forall s\in\mathcal{S}_{0}=\mathbb{C}_{+}.$ (32)
These yield
$|\Phi(s)|<\max_{\omega\in\mathbb{R}}|\Phi(j\omega)|\leq 1,\ \forall s\in\mathbb{C}_{+},$ (33)
which implies $\Phi(s)\neq 1,\ \forall s\in\mathbb{C}_{+}$, and $1-\Phi(s)$ has no roots in $\mathbb{C}_{+}$. Therefore, the open-loop transfer function (3) has no poles in the open right-half plane.
Appendix C Proof of Theorem 0
Let a simply connected region $\mathcal{C}_{\varepsilon}\subset\mathbb{C}_{+}$ and simple closed contour $\mathcal{D}_{\varepsilon,\delta}\subset\mathcal{C}_{\varepsilon}$ be
$\mathcal{C}_{\varepsilon}\coloneqq\{s\in\mathbb{C}\mid\mathrm{Re}[s]>\varepsilon>0\}$ (34a)
$\mathcal{D}_{\varepsilon,\delta}\coloneqq\{\delta+jr\mid-R_{\varepsilon,\delta}\leq r\leq R_{\varepsilon,\delta}\}\cup\mathcal{A}_{\varepsilon,\delta}$ (34b)
$\mathcal{A}_{\varepsilon,\delta}\coloneqq\{\delta+R_{\varepsilon,\delta}e^{j\theta}\mid\theta\in\Theta\}$ (34c)
$\Theta\coloneqq(-\pi/2,\pi/2),\ R_{\varepsilon,\delta}\coloneqq 2\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\phi_{\varepsilon}/(\delta-\varepsilon)$ (34d)
$\phi_{\varepsilon}\coloneqq 1/(1-\max_{\mathrm{Re}[s]=\varepsilon}|\Phi(s)|)\in[1,\infty)$ (34e)
$\delta\in\Delta\coloneqq\{\delta\in(\varepsilon,1/2]\mid R_{\varepsilon,\delta}>2(\delta+\omega_{\mathrm{b}})\}$ (34f)
using the linear-phase low-pass filter $\Phi(s)$ in (14a). On the region $\mathcal{A}_{\varepsilon,\delta}\subset\mathcal{S}_{\delta}$, the non-constant filter $\Phi(s)$ satisfies
$|\Phi(s)|<\max_{\mathrm{Re}[s]=\delta}|\Phi(s)|<\max_{\mathrm{Re}[s]=\varepsilon}|\Phi(s)|<1$ (35a)
${1}/(1-|\Phi(s)|)<\phi_{\varepsilon}<\infty$ (35b)
based on Lemma 4. Given $\varepsilon$, $\phi_{\varepsilon}$ is a finite constant independent of $\delta$. Using the following relationships
$\left|\frac{1+\Phi(s)}{1-\Phi(s)}\right|\leq 2\phi_{\varepsilon},\ \forall s\in\mathcal{A}_{\varepsilon,\delta}.$ (36a)
$R_{\varepsilon,\delta}>2(\delta+\omega_{\mathrm{b}})\Rightarrow-(\delta+\omega_{\mathrm{b}})>-R_{\varepsilon,\delta}/2$
$\Rightarrow 1/(R_{\varepsilon,\delta}-\delta-\omega_{\mathrm{b}})<2/R_{\varepsilon,\delta},$ (36b)
the open-loop transfer function (3) is bounded as
$\sup_{s\in\mathcal{A}_{\varepsilon,\delta}}|\Gamma(\delta+R_{\varepsilon,\delta}e^{j\theta})|\leq\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\phi_{\varepsilon}}{R_{\varepsilon,\delta}-\delta-\omega_{\mathrm{b}}}$ (37)
$\leq 2\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\phi_{\varepsilon}/R_{\varepsilon,\delta}=\delta-\varepsilon<1/2.$ (38)
Hence,
$\log(1+\Gamma(s))=\Gamma(s)$ $+\sum_{n=2}^{\infty}(-1)^{n+1}\frac{\Gamma(s)^{n}}{n}$ (39)
$|\log(1+\Gamma(s))-\Gamma(s)|$ $\leq\sum_{n=0}^{\infty}|\Gamma(s)|^{n}-1-|\Gamma(s)|$
$\leq\frac{|\Gamma(s)|^{2}}{1-|\Gamma(s)|}\leq 2|\Gamma(s)|^{2}.$ (40)
These and the estimation lemma give
$\left|\int_{\mathcal{A}_{\varepsilon,\delta}}\log(1+\Gamma(s))-\Gamma(s)ds\right|\leq 2\pi R_{\varepsilon,\delta}\sup_{s\in\mathcal{A}_{\varepsilon,\delta}}|\Gamma(s)|^{2}$ (41)
with the length $\pi R_{\varepsilon,\delta}$ of $\mathcal{A}_{\varepsilon,\delta}$. The limit of the upper bound equals zero as
$\lim_{\delta\to\varepsilon^{+}}$ $R_{\varepsilon,\delta}\sup_{s\in\mathcal{A}_{\varepsilon,\delta}}|\Gamma(s)|^{2}\leq\lim_{\delta\to\varepsilon^{+}}R_{\varepsilon,\delta}\left(\frac{2\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\phi_{\varepsilon}}{R_{\varepsilon,\delta}}\right)^{2}$
$\leq\lim_{\delta\to\varepsilon^{+}}2\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\phi_{\varepsilon}(\delta-\varepsilon)=0.$ (42)
Thus, $\lim_{\delta\to\varepsilon^{+}}|\int_{\mathcal{A}_{\varepsilon,\delta}}\log(1+\Gamma(s))-\Gamma(s)ds|=0$, and
$\lim_{\delta\to\varepsilon^{+}}\int_{\mathcal{A}_{\varepsilon,\delta}}\log(1+\Gamma(s))ds=\lim_{\delta\to\varepsilon^{+}}\int_{\mathcal{A}_{\varepsilon,\delta}}\Gamma(s)ds.$ (43)
In $\mathbb{C}_{+}$, the sensitivity function $S(s)=1/(1+\Gamma(s))$ has no zeros owing to Lemma 5 and no poles because the QDOB is stable according to Muramatsu (2025) with (33). Hence, $S(s)$ is analytic in $\mathbb{C}_{+}$. For the analytic function $S(s)$ that has no zeros in the simply connected region $\mathcal{C}_{\varepsilon}$, there exists a single-valued analytic branch of the logarithm of $S(s)$ in $\mathcal{D}_{\varepsilon,\delta}$ according to Ahlfors (1979). Let $\log(S(s))$ denote this single-valued analytic branch such that $\log(S(s))=\ln|S(s)|+j\angle S(s)$. We apply the Cauchy integral theorem to the branch $\log(S(s))$ on $\mathcal{D}_{\varepsilon,\delta}$ traversed counterclockwise:
$\oint_{\mathcal{D}_{\varepsilon,\delta}}\log(S(s))ds$ $=0.$ (44)
Then, (44) can be transformed as follows
$\int_{\delta+jR_{\varepsilon,\delta}}^{\delta-jR_{\varepsilon,\delta}}\log(S(s))ds+\int_{\mathcal{A}_{\varepsilon,\delta}}\log(S(s))ds=0$ (45a)
$j\int_{-R_{\varepsilon,\delta}}^{R_{\varepsilon,\delta}}\log(S(\delta+j\omega))d\omega=\int_{\mathcal{A}_{\varepsilon,\delta}}\log(S(s))ds$ (45b)
$\int_{-R_{\varepsilon,\delta}}^{R_{\varepsilon,\delta}}\ln|S(\delta+j\omega)|d\omega=\mathrm{Im}\left[\int_{\mathcal{A}_{\varepsilon,\delta}}\log(S(s))ds\right].$ (45c)
The Bode-like sensitivity integral is calculated as
$\lim_{\varepsilon\to 0^{+}}\int_{0}^{\infty}\ln|S(\varepsilon+j\omega)|d\omega=\lim_{\varepsilon\to 0^{+}}\frac{1}{2}\int_{-\infty}^{\infty}\ln|S(\varepsilon+j\omega)|d\omega$
$=\frac{1}{2}\lim_{\varepsilon\to 0^{+}}\lim_{\delta\to\varepsilon^{+}}\int_{-R_{\varepsilon,\delta}}^{R_{\varepsilon,\delta}}\ln|S(\delta+j\omega)|d\omega$
$=\frac{1}{2}\lim_{\varepsilon\to 0^{+}}\lim_{\delta\to\varepsilon^{+}}\mathrm{Im}\left[\int_{\mathcal{A}_{\varepsilon,\delta}}\log(S(s))ds\right]$
$=-\frac{1}{2}\lim_{\varepsilon\to 0^{+}}\lim_{\delta\to\varepsilon^{+}}\mathrm{Im}\left[\int_{\mathcal{A}_{\varepsilon,\delta}}\log(1+\Gamma(s))ds\right].$ (46)
From (43),
$\lim_{\delta\to\varepsilon^{+}}\int_{\mathcal{A}_{\varepsilon,\delta}}\log(1+\Gamma(s))ds=\lim_{\delta\to\varepsilon^{+}}\int_{\mathcal{A}_{\varepsilon,\delta}}\Gamma(s)ds$
$=j\lim_{\delta\to\varepsilon^{+}}\int_{-\pi/2}^{\pi/2}\Gamma(\delta+R_{\varepsilon,\delta}e^{j\theta})R_{\varepsilon,\delta}e^{j\theta}d\theta.$ (47)
We apply the dominated convergence theorem. Because
$\lim_{\delta\to\varepsilon^{+}}R_{\varepsilon,\delta}=\infty$ (48a)
$\lim_{\delta\to\varepsilon^{+}}|\Phi(\delta+R_{\varepsilon,\delta}e^{j\theta})|$
$\hskip 10.00002pt\leq\lim_{\delta\to\varepsilon^{+}}\prod_{i=1}^{l}\sum_{n=-N}^{N}\left|\frac{\alpha_{i}(n)}{\gamma_{i}}\right|e^{((n-N)\bar{U}_{i}-\kappa)TR_{\varepsilon,\delta}\cos(\theta)}$
$\hskip 10.00002pt=0$ (48b)
$\lim_{\delta\to\varepsilon^{+}}\Phi(\delta+R_{\varepsilon,\delta}e^{j\theta})=0,\ \forall\theta\in\Theta,$ (48c)
the function $\Gamma(\delta+R_{\varepsilon,\delta}e^{j\theta})R_{\varepsilon,\delta}e^{j\theta}$ converges pointwise
$\lim_{\delta\to\varepsilon^{+}}\Gamma(\delta+R_{\varepsilon,\delta}e^{j\theta})R_{\varepsilon,\delta}e^{j\theta}$
$=\lim_{\delta\to\varepsilon^{+}}\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}{2}\frac{1+\Phi(\delta+R_{\varepsilon,\delta}e^{j\theta})}{1-\Phi(\delta+R_{\varepsilon,\delta}e^{j\theta})}\frac{R_{\varepsilon,\delta}e^{j\theta}}{\delta+R_{\varepsilon,\delta}e^{j\theta}+\omega_{\mathrm{b}}}$
$={\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}/{2}.$ (49)
Additionally,
$|\Gamma(\delta+R_{\varepsilon,\delta}e^{j\theta})R_{\varepsilon,\delta}e^{j\theta}|\leq 2\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\phi_{\varepsilon},$ (50)
where the upper bound is an integrable function independent of $\delta$. Thus, the limit and integral can be interchanged as
$\lim_{\delta\to\varepsilon^{+}}\int_{-\pi/2}^{\pi/2}\Gamma(\delta+R_{\varepsilon,\delta}e^{j\theta})R_{\varepsilon,\delta}e^{j\theta}d\theta$
$=\int_{-\pi/2}^{\pi/2}\lim_{\delta\to\varepsilon^{+}}\Gamma(\delta+R_{\varepsilon,\delta}e^{j\theta})R_{\varepsilon,\delta}e^{j\theta}d\theta$
$=\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}{2}\int_{-\pi/2}^{\pi/2}d\theta=\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\pi}{2}.$ (51)
From (C), (C), and (C), the Bode-like sensitivity integral for the QDOB is obtained as
$\lim_{\varepsilon\to 0^{+}}\int_{0}^{\infty}\ln|S(\varepsilon+j\omega)|d\omega=-\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\pi}{4}.$ (52)
Appendix D Proof of Theorem 0
Let a simply connected region $\mathcal{C}_{\varepsilon}\subset\mathbb{C}$ be
$\mathcal{C}_{\varepsilon}=\{\tilde{z}\in\mathbb{C}\mid|\tilde{z}|<\varepsilon^{-1}<1\},\ \tilde{z}\coloneqq z^{-1}.$ (53)
Suppose the linear-phase low-pass filter $\tilde{\Phi}(z)$ in (17), open-loop transfer function $\tilde{\Gamma}(z)$ in (16), and sensitivity function $\tilde{S}(z)$. We substitute $z=\tilde{z}^{-1}$ for the functions, as follows
[…]
$)\displaystyle\log(\hat{S}(\tilde{z}$)) $=\log(\hat{S}(0))+a_{1}\tilde{z}+a_{2}\tilde{z}^{2}+\cdots$ (55)
with $a_{1},\ a_{2},\ \ldots\in\mathbb{C}$. Accordingly,
$\log(\tilde{S}(z))=\log(\hat{S}(0))+a_{1}z^{-1}+a_{2}z^{-2}+\cdots$ (56)
Using this series, the contour integral of $\log(\tilde{S}(z))z^{-1}$ along $|z|=\varepsilon$ is
$\lim_{\varepsilon\to 1^{+}}$ $\oint_{|z|=\varepsilon}\log(\tilde{S}(z))z^{-1}dz$
$=\lim_{\varepsilon\to 1^{+}}\oint_{|z|=\varepsilon}\log(\hat{S}(0))z^{-1}+a_{1}z^{-2}+\cdots dz$
$=j\lim_{\varepsilon\to 1^{+}}\int_{0}^{2\pi}\log(\hat{S}(0))+a_{1}\varepsilon^{-1}e^{-j\theta}+\cdots d\theta$
$=j2\pi\log(\hat{S}(0)).$ (57)
Meanwhile, the contour integral becomes
$\lim_{\varepsilon\to 1^{+}}$ $\oint_{|z|=\varepsilon}\log(\tilde{S}(z))z^{-1}dz$
$=\lim_{\varepsilon\to 1^{+}}\int_{0}^{2\pi}j\ln|\tilde{S}(\varepsilon e^{j\Omega})|-\angle\tilde{S}(\varepsilon e^{j\Omega})d\Omega.$ (58)
From the imaginary parts of (D) and (D), the Bode-like sensitivity integral for the discrete-time representation is
$\lim_{\varepsilon\to 1^{+}}\int_{0}^{2\pi}\ln|\tilde{S}(\varepsilon e^{j\Omega})|d\Omega=2\pi\mathrm{Re}[\log(\hat{S}(0))]$
$=2\pi\ln(2+2\omega_{\mathrm{b}}T)-2\pi\ln(2+2\omega_{\mathrm{b}}T+\omega_{\mathrm{b}}\omega_{\mathrm{c}}LT).$ (59)